# New Convergence Analysis of GMRES with

### v1

@@ Abstract
New convergence bounds are presented for weighted, preconditioned, and deflated GMRES for the solution of large, sparse, nonsymmetric linear systems, where it is assumed that the symmetric part of the coefficient matrix is positive definite. The new bounds are sufficiently explicit to indicate how to choose the preconditioner and the deflation space to accelerate the convergence. One such choice of deflating space is presented, and numerical experiments illustrate the effectiveness of such space.
@@ 1 Introduction
where $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and $\mathbf{A}$ is a large sparse nonsingular matrix. Particular emphasis will be on the cases where $\mathbf{A}$ has a positive definite Hermitian part. We refer to such matrices $\mathbf{A}$ as positive definite matrices.
@@
Standard references for preconditioning include [2, 18, 20]; for weighted GMRES [8, 10, 14]; and for deflation [4, 6, 9, 12], and more recently [11]. Here, we denote the weighted preconditioned and deflated GMRES algorithm as WPD-GMRES, and corresponds to the case where all three acceleration tools are used.
Our objective in this paper is to propose a new convergence bound for WPD-GMRES that is sufficiently explicit to indicate how to choose the preconditioner, weight matrix, and especially deflation spaces. The new results generalize those in [22] where deflation was not considered. Here also, special emphasis is on Hermitian preconditioning and on applying WPD-GMRES in the preconditioner norm, an idea already present in [5, 25].
In particular, in Section 6, we present a result explicitly giving conditions on the preconditioner and the deflation spaces so as to assure fast convergence. Then, in Section 7, we propose a new deflation space which is inspired by this new bound. Numerical experiments in Section 8 illustrate the new results with the choice of new space, and show its effectiveness.
@@ 2 Preliminaries
GMRES for singular systems is studied in [3] and [15] (where GCR is also considered). A very useful result is recalled in Theorem 1 with a straightforward generalization to weighted GMRES. The proof of [3, Theorem 2.6] applies here with the change of inner product.
@@
Suppose that $\operatorname{range}(\mathbf{A})\cap\operatorname{ker}(\mathbf{A})=\{0\}$. If $\mathbf{A}\mathbf{x}=\mathbf{b}$ is consistent, i.e., if it admits a solution, then weighted GMRES determines a solution without breakdown at some step and breaks down at the next step through degeneracy of the Krylov subspace.
The takeaway from the theorem is that for consistent linear systems, under the condition that $\operatorname{range}(\mathbf{A})\cap\operatorname{ker}(\mathbf{A})=\{\mathbf{0}\}$, we can proceed through the iterations in the same manner as with nonsingular systems. In exact arithmetic, the characterization (2) of the iterate $\mathbf{x}_{i}$ remains valid until the algorithm breaks down, at which point the exact solution has been found.
We continue in this preliminaries section by reviewing properties of the generalized eigenvalue problem we use for our new deflation space.
@@
Let us assume that $\mathbf{M}$ and $\mathbf{N}$ are two order $n$ matrices with the further assumption that $\mathbf{M}$ is hpd and $\mathbf{N}$ is skew-Hermitian. Consider the generalized eigenvalue problem for matrix pencil $(\mathbf{N},\mathbf{M})$: find $\lambda_{j}\in\mathbb{C}$ and $\mathbf{z}^{(j)}\in\mathbb{C}^{n}\setminus\{\mathbf{0}\}$ such that
@@
where $\mathbf{M}^{1/2}$ denotes the matrix square root of $\mathbf{M}$ [^1] By [17, theorem 7.2.6, page 439], $\mathbf{M}^{1/2}$ is well defined as the unique Hermitian positive semi-definite matrix such that $(\mathbf{M}^{1/2})^{2}=\mathbf{M}$ and moreover $\mathbf{M}^{1/2}$ is positive-definite because $\mathbf{M}$ is positive-definite.
Matrix $\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2}$ is skew-symmetric: $(\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2})^{*}=(\mathbf{M}^{-1/2})^{*}\mathbf{N}^{*}(\mathbf{M}^{-1/2})^{*}=-\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2}$. Consequently $\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2}$ is normal and the spectral theorem states that it is unitarily diagonalizable:
@@
We prove a few straightforward properties of the eigenpairs. If $(\lambda_{k},\mathbf{z}^{(k)})$ denotes an eigenpair of the generalized eigenvalue problem (4) then
@@
3 Weighted and Deflated GMRES
@@
and setting $\mathbf{Q}_{D}\mathbf{x}_{*}=\mathbf{Q}_{D}\tilde{}\mathbf{x}$. This is justified by [11, Lemma 3.2] or by the following one-line proof,
@@
In those references, and in this paper, it is implicitly assumed that the number of columns $m$ of $\mathbf{Y}$ and $\mathbf{Z}$ is not too large so that the solution of the solution of a linear system with the $m\times m$ coefficient matrix $\mathbf{Y}^{*}\mathbf{A}\mathbf{Z}$ is not too expensive. Such solutions are needed when computing $(\mathbf{{I}}-\mathbf{Q}_{D})\mathbf{x}_{*}$ as in (6), and at every application of $\mathbf{P}_{D}$ and $\mathbf{Q}_{D}$.
@@ 4 Weighted and Deflated right-preconditioned GMRES
Equation (8) is a consistent linear system with a singular coefficient matrix $\mathbf{P}_{D}\mathbf{A}\mathbf{H}$. By Theorem 1 (see also [11, Theorem 3.4]), weighted and deflated preconditioned GMRES converges for every starting vector if
@@
This is the same condition as in [11, Theorem 3.5] where left preconditioning is considered. Indeed, the Krylov subspaces ${\mathcal{K}_{i}}(\mathbf{P}_{D}\mathbf{A}\mathbf{H},\mathbf{P}_{D}\mathbf{b}-\mathbf{P}_{D}\mathbf{A}\mathbf{H}\mathbf{u}_{0})$ with $\mathbf{u}_{0}=\mathbf{H}\mathbf{x}_{0}$ and ${\mathcal{K}_{i}}(\mathbf{H}\mathbf{P}_{D}\mathbf{A},\mathbf{H}\mathbf{P}_{D}\mathbf{b}-\mathbf{H}\mathbf{P}_{D}\mathbf{A}\mathbf{x}_{0})$ stop growing at the same iteration, i.e., the coefficient matrices have the same grades in the sense of [18, Section 6.2]. Indeed, ${\mathcal{K}_{i}}(\mathbf{H}\mathbf{P}_{D}\mathbf{A},\mathbf{H}\mathbf{P}_{D}\mathbf{b}-\mathbf{H}\mathbf{P}_{D}\mathbf{A}\mathbf{x}_{0})=\mathbf{H}{\mathcal{K}_{i}}(\mathbf{P}_{D}\mathbf{A}\mathbf{H},\mathbf{P}_{D}\mathbf{b}-\mathbf{P}_{D}\mathbf{A}\mathbf{H}\mathbf{u}_{0})$ with $\mathbf{x}_{0}=\mathbf{H}\mathbf{u}_{0}$.
@@
5 Convergence of WPD-GMRES
@@
Proof.
By Theorem 2, there is no breakdown until convergence has been achieved. So at iteration $i$ of weighted GMRES applied to (8) it holds that
@@
The result follows by dividing by $\|\mathbf{r}_{i-1}\|_{\mathbf{W}}^{2}$ and recalling that $\mathbf{r}_{i-1}\in\operatorname{range}(\mathbf{P}_{D})$. ∎
@@
The convergence bound in Theorem 3 is pessimistic for GMRES. Indeed, it is derived from $\|\mathbf{r}_{i}\|_{\mathbf{W}}\leq\operatorname{min}\left\{\|\mathbf{P}_{D}\mathbf{b}-\mathbf{P}_{D}\mathbf{A}\mathbf{x}\|_{\mathbf{W}};\,{\mathbf{x}\in\mathbf{x}_{i-1}+\operatorname{span}(\mathbf{H}\mathbf{r}_{i-1})}\right\}$ where the global minimization property of GMRES has been deteriorated to minimizing over a one-dimensional space. For this reason, the bound in Theorem 3 holds also for all restarted and truncated versions of GMRES and even for the minimal residual algorithm. The remark carries over to all convergence results in the article since they are essentially bounds for $\theta(\mathbf{A},\mathbf{H},\mathbf{W},\mathbf{Y},\mathbf{Z})$.
@@
6 Hpd preconditioning for $\mathbf{A}$ positive definite
@@
From here, two cases are considered that differ by the constraint imposed on the deflation spaces. In each case, the objective is to make explicit a condition that must be satisfied by $\mathbf{H}$, $\mathbf{Y}$ and $\mathbf{Z}$ in order to ensure fast convergence. The results are summed up in the next theorem
@@
    $\theta(\mathbf{A},\mathbf{H},\mathbf{H},\mathbf{Y},\mathbf{Z})\geq\frac{\lambda_{\min}(\mathbf{H}\mathbf{M})}{\lambda_{\max}(\mathbf{H}\mathbf{M})}\times\operatorname{inf}\limits_{\mathbf{y}\in\ker(\mathbf{Z}^{*}\mathbf{A}^{*}\mathbf{A}^{-1})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\mathbf{M}^{-1}}\mathbf{y}\rangle}\cdot$
@@
    This case corresponds to Lemma 3 where it is proved that only the condition $\operatorname{ker}(\mathbf{Y}^{*})\cap\operatorname{range}(\mathbf{H}^{-1}\mathbf{Z})=\{\mathbf{0}\}$ is necessary in order to ensure that WPD-GMRES does not break down. Here, that condition is equivalent to $\operatorname{ker}(\mathbf{Z}^{*}\mathbf{A}^{*})\cap\operatorname{range}(\mathbf{Z})=\{\mathbf{0}\}$. Let’s take a vector $\mathbf{Z}\mathbf{z}$ in that intersection: $\mathbf{Z}^{*}\mathbf{A}^{*}\mathbf{Z}\mathbf{z}=0$ implies that
@@
    where $\mathbb{I}$ stands for the imaginary axis. The positive definiteness of $\mathbf{A}$ allows us to conclude that $\mathbf{z}=\mathbf{0}$ and that $\operatorname{ker}(\mathbf{Z}^{*}\mathbf{A}^{*})\cap\operatorname{range}(\mathbf{Z})=\{\mathbf{0}\}$.
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{H}\mathbf{P}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{H}\mathbf{P}_{D})\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{y},\mathbf{H}\mathbf{y}\rangle}$
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{H}\mathbf{P}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}\mathbb{K}^{n}\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{y},\mathbf{H}\mathbf{y}\rangle}$
@@
    $\theta(\mathbf{A},\mathbf{H},\mathbf{H},\tilde{}\mathbf{Y},\mathbf{Z})$ $=\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}\mathbf{y},\mathbf{y}\rangle|^{2}}{\langle\mathbf{H}\mathbf{A}\mathbf{y},\mathbf{A}\mathbf{y}\rangle\langle\mathbf{H}^{-1}\mathbf{y},\mathbf{y}\rangle}$
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{H}\mathbf{y},\mathbf{y}\rangle}$
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\mathbb{K}^{n}\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{H}\mathbf{y},\mathbf{y}\rangle}$

### v2

@@ Abstract
New convergence bounds are presented for weighted, preconditioned, and deflated GMRES for the solution of large, sparse, non-Hermitian linear systems. These bounds are given for the case when the Hermitian part of the coefficient matrix is positive definite, the preconditioner is Hermitian positive definite, and the weight is equal to the preconditioner. The new bounds are a novel contribution in and of themselves. In addition, they are sufficiently explicit to indicate how to choose the preconditioner and the deflation space to accelerate the convergence. One such choice of deflating space is presented, and numerical experiments illustrate the effectiveness of such space.
@@ 1 Introduction
where $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and $\mathbf{A}$ is a large sparse nonsingular matrix. Particular emphasis will be on the cases where $\mathbf{A}$ has a positive definite Hermitian (hpd) part. We refer to such matrices $\mathbf{A}$ as positive definite (pd) matrices.
@@
Standard references for preconditioning include [2, 21, 24]; for weighted GMRES [10, 12, 16]; and for deflation [4, 6, 11, 14], and more recently [13]. See also [33], where a weighted and deflated version of GMRES is applied to Sylvester matrix equations. There are some equivalences between weighting and preconditioning, see, e.g., [20, Chapter 4], but here we keep the weight and the preconditioner separate for further flexibility and generality. We denote the weighted preconditioned and deflated GMRES algorithm as WPD-GMRES, and it corresponds to the case where all three acceleration tools are used.
Our objective in this paper is to propose a new convergence bound for WPD-GMRES that is sufficiently explicit to indicate how to choose the preconditioner, weight matrix, and especially deflation spaces. The new results generalize those in [26] where deflation was not considered. Here also, special emphasis is on Hermitian preconditioning and on applying WPD-GMRES in the preconditioner norm, as was done, e.g., in [5, 29]. For the cases we consider, the new bounds improve upon the more general bound of Elman [8, 9].
In particular, in Section 5, we present a result explicitly giving conditions on the preconditioner and the deflation spaces so as to assure fast convergence. Then, in Section 6, we propose a new deflation space which is inspired by this new bound. Numerical experiments in Section 7 illustrate the new results with the choice of the new space, and show its effectiveness. In fact, in one of the example problems we present, GMRES fails to converge, while a deflation space with dimension less than 10% of the size of the problems, allows GMRES to converge.
@@
In summary, our contribution consists of proving a new convergence bound for weighted preconditioned GMRES, for the special case that the (right) preconditioneer is the same as the weight, and $A$ is positive definite; and to provide a special deflating space which allows us to quantify this bound explicitly.
@@ 2 Preliminaries
GMRES for singular systems is studied, e.g, in [3] and [17] (where GCR is also considered). A very useful result is recalled in Theorem 1 with a straightforward generalization to weighted GMRES. The proof of [3, Theorem 2.6] applies here with the change of inner product; see also [14].
@@
Suppose that $\operatorname{range}(\mathbf{A})\cap\operatorname{ker}(\mathbf{A})=\{0\}$. If $\mathbf{A}\mathbf{x}=\mathbf{b}$ is consistent, i.e., if it admits a solution, then, in exact arithmetic, weighted GMRES determines a solution without breakdown at some step and breaks down at the next step through degeneracy of the Krylov subspace.
The takeaway from the theorem is that for consistent linear systems, under the condition that $\operatorname{range}(\mathbf{A})\cap\operatorname{ker}(\mathbf{A})=\{\mathbf{0}\}$, we can proceed through the iterations in the same manner as with nonsingular systems. In exact arithmetic, the characterization (2) of the iterate $\mathbf{x}_{i}$ remains valid in the singular case, until the algorithm breaks down, at which point the exact solution has been found.
We continue in this preliminaries section by reviewing properties of the generalized eigenvalue problem we use for our new deflation space. For completeness we give the proof of the following result.
@@
Let us assume that $\mathbf{M}$ and $\mathbf{N}$ are two $n\times n$ matrices with the further assumption that $\mathbf{M}$ is hpd and $\mathbf{N}$ is skew-Hermitian. Consider the generalized eigenvalue problem for matrix pencil $(\mathbf{N},\mathbf{M})$: find $\lambda_{j}\in\mathbb{C}$ and $\mathbf{z}^{(j)}\in\mathbb{C}^{n}\setminus\{\mathbf{0}\}$ such that
@@
where $\mathbf{M}^{1/2}$ denotes the matrix square root of $\mathbf{M}$ [^1] By [19, theorem 7.2.6, page 439], $\mathbf{M}^{1/2}$ is well defined as the unique Hermitian positive semi-definite matrix such that $(\mathbf{M}^{1/2})^{2}=\mathbf{M}$ and moreover $\mathbf{M}^{1/2}$ is positive-definite because $\mathbf{M}$ is positive-definite.
Matrix $\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2}$ is skew-Hermitian: $(\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2})^{*}=(\mathbf{M}^{-1/2})^{*}\mathbf{N}^{*}(\mathbf{M}^{-1/2})^{*}=-\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2}$. Consequently $\mathbf{M}^{-1/2}\mathbf{N}\mathbf{M}^{-1/2}$ is normal and the spectral theorem states that it is unitarily diagonalizable:
@@
We prove a few straightforward properties of the eigenpairs, which we will use later in the paper. If $(\lambda_{k},\mathbf{z}^{(k)})$ denotes an eigenpair of the generalized eigenvalue problem (4) then
@@
2.1 Weighted and Deflated GMRES
@@
and setting $\mathbf{Q}_{D}\mathbf{x}_{*}=\mathbf{Q}_{D}\tilde{}\mathbf{x}$. This is justified by the following one-line proof (which is essentially [13, Lemma 3.2]),
@@
In those references and in this paper, it is implicitly assumed that the number of columns $m$ of $\mathbf{Y}$ and $\mathbf{Z}$ is not too large so that the solution of the solution of a linear system with the $m\times m$ coefficient matrix $\mathbf{Y}^{*}\mathbf{A}\mathbf{Z}$ is not too expensive. Such solutions are needed when computing $(\mathbf{{I}}-\mathbf{Q}_{D})\mathbf{x}_{*}$ as in (6), and at every application of $\mathbf{P}_{D}$ and $\mathbf{Q}_{D}$.
@@ 3 Weighted and Deflated right-preconditioned GMRES
Equation (8) is a consistent linear system with a singular coefficient matrix $\mathbf{P}_{D}\mathbf{A}\mathbf{H}$. By Theorem 1 (see also [13, Theorem 3.4]), weighted and deflated preconditioned GMRES converges for every starting vector if
@@
This is the same condition as in [13, Theorem 3.5] where left preconditioning is considered. Indeed, the Krylov subspaces ${\mathcal{K}_{i}}(\mathbf{P}_{D}\mathbf{A}\mathbf{H},\mathbf{P}_{D}\mathbf{b}-\mathbf{P}_{D}\mathbf{A}\mathbf{H}\mathbf{u}_{0})$ with $\mathbf{u}_{0}=\mathbf{H}\mathbf{x}_{0}$ and ${\mathcal{K}_{i}}(\mathbf{H}\mathbf{P}_{D}\mathbf{A},\mathbf{H}\mathbf{P}_{D}\mathbf{b}-\mathbf{H}\mathbf{P}_{D}\mathbf{A}\mathbf{x}_{0})$ stop growing at the same iteration, i.e., the coefficient matrices have the same grades in the sense of [21, Section 6.2]. Indeed, ${\mathcal{K}_{i}}(\mathbf{H}\mathbf{P}_{D}\mathbf{A},\mathbf{H}\mathbf{P}_{D}\mathbf{b}-\mathbf{H}\mathbf{P}_{D}\mathbf{A}\mathbf{x}_{0})=\mathbf{H}{\mathcal{K}_{i}}(\mathbf{P}_{D}\mathbf{A}\mathbf{H},\mathbf{P}_{D}\mathbf{b}-\mathbf{P}_{D}\mathbf{A}\mathbf{H}\mathbf{u}_{0})$ with $\mathbf{x}_{0}=\mathbf{H}\mathbf{u}_{0}$.
@@
4 General Convergence of WPD-GMRES
The following result extends [26, Theorem 3] to the case where we have a deflation space.
@@ 4 General Convergence of WPD-GMRES
Note that by Cauchy-Schwarz inequality for the $\mathbf{W}$-norm, we have that $\theta(\mathbf{A},\mathbf{H},\mathbf{W},\mathbf{Y},\mathbf{Z})<1$.
Proof of Theorem 3. By Theorem 2, there is no breakdown until convergence has been achieved. So at iteration $i$ of weighted GMRES applied to (8) it holds that
@@
The result follows by dividing by $\|\mathbf{r}_{i-1}\|_{\mathbf{W}}^{2}$ and recalling that $\mathbf{r}_{i-1}\in\operatorname{range}(\mathbf{P}_{D})$.
@@
The convergence bound in Theorem 3 is pessimistic for GMRES. Indeed, it is derived from $\|\mathbf{r}_{i}\|_{\mathbf{W}}\leq\operatorname{min}\left\{\|\mathbf{P}_{D}\mathbf{b}-\mathbf{P}_{D}\mathbf{A}\mathbf{x}\|_{\mathbf{W}};\,{\mathbf{x}\in\mathbf{x}_{i-1}+\operatorname{span}(\mathbf{H}\mathbf{r}_{i-1})}\right\}$ where the global minimization property of GMRES has been bounded by minimizing over a one-dimensional space. For this reason, the bound in Theorem 3 holds also for all restarted and truncated versions of GMRES and even for the minimal residual algorithm. The remark carries over to all convergence results in the article since they are essentially bounds for $\theta(\mathbf{A},\mathbf{H},\mathbf{W},\mathbf{Y},\mathbf{Z})$.
@@
5 The case of hpd preconditioning for $\mathbf{A}$ positive definite
@@
From here, two cases are considered that differ by the constraint imposed on the deflation spaces. In each case, the objective is to make explicit a condition that must be satisfied by $\mathbf{H}$, $\mathbf{Y}$ and $\mathbf{Z}$ in order to ensure fast convergence. The results are summed up in the next theorem. The following is our main result which extends [26, Theorem 6] to the case where we have a deflation space.
@@
    $\theta(\mathbf{A},\mathbf{H},\mathbf{H},\mathbf{H}\mathbf{A}\mathbf{Z},\mathbf{Z})\geq\frac{\lambda_{\min}(\mathbf{H}\mathbf{M})}{\lambda_{\max}(\mathbf{H}\mathbf{M})}\times\operatorname{inf}\limits_{\mathbf{y}\in\ker(\mathbf{Z}^{*}\mathbf{A}^{*}\mathbf{A}^{-1})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\mathbf{M}^{-1}}\mathbf{y}\rangle}\cdot$
@@
    This case corresponds to Lemma 3 where it is proved that only the condition $\operatorname{ker}(\mathbf{Y}^{*})\cap\operatorname{range}(\mathbf{H}^{-1}\mathbf{Z})=\{\mathbf{0}\}$ is necessary in order to ensure that WPD-GMRES does not break down. Here, that condition is equivalent to $\operatorname{ker}(\mathbf{Z}^{*}\mathbf{A}^{*})\cap\operatorname{range}(\mathbf{Z})=\{\mathbf{0}\}$. To prove this, consider a vector $\mathbf{Z}\mathbf{z}$ in that intersection: $\mathbf{Z}^{*}\mathbf{A}^{*}\mathbf{Z}\mathbf{z}=0$ implies that
@@
    where $\mathbb{I}$ stands for the imaginary axis. Thus, $\langle\mathbf{Z}^{*}\mathbf{M}\mathbf{Z}\mathbf{z},\mathbf{z}\rangle=0$ and since $M$ is positive definite, $\mathbf{Z}\mathbf{z}=0$, and this allows us to conclude that $\mathbf{z}=\mathbf{0}$ and that $\operatorname{ker}(\mathbf{Z}^{*}\mathbf{A}^{*})\cap\operatorname{range}(\mathbf{Z})=\{\mathbf{0}\}$.
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{H}\mathbf{P}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{H}\mathbf{P}_{D})\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{y},\mathbf{H}\mathbf{y}\rangle}$
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{H}\mathbf{P}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}\mathbb{K}^{n}\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{y},\mathbf{H}\mathbf{y}\rangle}$
@@
    $\theta(\mathbf{A},\mathbf{H},\mathbf{H},\mathbf{Y},\mathbf{Z})$ $=\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}\mathbf{y},\mathbf{y}\rangle|^{2}}{\langle\mathbf{H}\mathbf{A}\mathbf{y},\mathbf{A}\mathbf{y}\rangle\langle\mathbf{H}^{-1}\mathbf{y},\mathbf{y}\rangle}$
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{H}\mathbf{y},\mathbf{y}\rangle}$
@@
    $\geq\operatorname{inf}\limits_{\mathbf{y}\in\operatorname{range}(\mathbf{A}\mathbf{Q}_{D})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}\times\operatorname{inf}\limits_{\mathbf{y}\in\mathbb{K}^{n}\setminus\{\mathbf{0}\}}\frac{\langle{\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}\mathbf{y},\mathbf{y}\rangle}{\langle\mathbf{H}\mathbf{y},\mathbf{y}\rangle}$